Solve for a: 2log5(3a+1)=log5(4)+2log5(2a-1)
Found 2 solutions by stanbon, MathTherapy:
Answer by stanbon(75887)

Solve for a: 2log5(3a+1)=log5(4)+2log5(2a-1)
-------------------
log5(3a+1)^2 - log5(2a-1)^2 = log(4)/log(5)
------
log5[(3a+1)^2/(2a-1)^2] = 0.8614
------
2log5[(3a+1)/(2a-1)] = 0.8614
-----
log5[(3a+1)/(2a-1)] = 0.4307
----
3a+1 = 5^0.4307*(2a-1)
3a+1 = 2(2a-1)
3a+1 = 4a-2
a = 3
---------------
Cheers,
Stan H.
----------------

Answer by MathTherapy(10809)

Solve for a: 2log5(3a+1)=log5(4)+2log5(2a-1)